OpenStudy (anonymous):

A chair of weight 115 N lies atop a horizontal floor; the floor is not frictionless. You push on the chair with a force of F = 38.0 N directed at an angle of 43.0^\circ below the horizontal and the chair slides along the floor.Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.

OpenStudy (anonymous):

Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.

OpenStudy (y2o2):

what is n ?!

OpenStudy (anonymous):

normal force

OpenStudy (anonymous):

The normal force must oppose all forces in the vertical direction because the chair doesn't move into the floor! \[0 = mg - F \sin(\theta)\]